Status Report

NASA ISS On-Orbit Status 24 August 2008

By SpaceRef Editor
August 24, 2008
Filed under , , ,
NASA ISS On-Orbit Status 24 August 2008

All ISS systems continue to function nominally, except those noted previously or below. Sunday – rest day for CDR Volkov, FE-1 Kononenko & FE-2 Chamitoff. Ahead: Week 19 of Increment 17.

Sleep/Wake Cycle Change: Crew sleep begins an hour earlier tonight (4:30pm EDT) to accommodate tomorrow’s TORU testing over RGS (Russian Groundsite) with a one-hour earlier wake-up (1:00am). Tomorrow evening, sleep will begin also earlier, at 4:30pm, and end at the regular time of 2:00am on Tuesday.

Gregory Chamitoff had three hours reserved for the regular weekly station cleaning in the USOS (US Segment) which wasn’t scheduled yesterday alongside the RS (Russian Segment) “uborka” housecleaning by his two crewmates.

Oleg Kononenko serviced the Russian BMP (Harmful Impurities Removal System), starting the "bake-out"-to-vacuum cycle on absorbent bed #2 of the regenerable dual-channel filtration system. The regen process will be terminated tonight at ~4:00pm EDT. Filter bed #1 was regenerated yesterday. [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP’s regeneration cycle, normally done every 20 days, is currently performed four times more frequently to remove any lingering Freon-218 from the cabin atmosphere (last time: 8/11&12).]

Sergey Volkov conducted the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the Service Module (SM), including the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP-Moscow. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The CDR also checked up on the Russian POTOK-150MK (150 micron) decontamination filter unit of the SM’s SOGS air revitalization subsystem, gathering weekly data on Total Operating Time & “On” durations for reporting to TsUP-Moscow.

Chamitoff completed the periodic relocation of the TEPC (Tissue Equivalent Proportional Counter), the primary radiation measurement tool in the ISS, from the SM starboard crew quarters (panel 449) to the JAXA JPM (JEM Pressurized Module) 1A5 location. [The FE-2 also photo documented the current state of the TEPC power/data cable, then covered the peeling Kapton tape on the degraded cable with fresh tape to contain any FOD (Foreign Object/Debris). TEPC had been relocated to the SM by Greg on 6/24.]

Working off the Russian discretionary “time permitting” task list, Kononenko performed the regular status check of the BIO-5 Rasteniya-2 ("Plants-2") experiment which researches growth and development of plants (peas) under spaceflight conditions in the Lada-13 greenhouse from IBMP (Institute of Bio-Medical Problems {Russian: IMBP}).

As time permitted, Oleg was also to conduct another run of the Russian DZZ-2 "Diatomeya" ocean observations program, using the NIKON-F5 DCS still camera with f85mm lens and the HDV camcorder from SM window #7 to record color field patterns in target water areas and current atmospheric conditions above them. [Uplinked target zones in the Indian Ocean were the coastal area of Oman and the area west of Australia, and in the Atlantic Ocean Newfoundland Island and the – Bay of Biscay.]

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-2), RED resistive exercise device (CDR, FE-1, FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1).

As generally every day now, today starting at ~9:00am and running until 3:00pm, the US CDRA (Carbon Dioxide Removal Assembly) was activated intermittently for two half-cycles to control ppCO2 levels. This configuration for the daily ops does not require connecting & disconnecting the ITCS (Internal Thermal Control System) cooling loop. [A forward plan is in work for cycling the CSV (CO2 Selector Valve) to prevent its sticking.]

Preparatory to ATV “Jules Verne” undocking on 9/5, TsUP-Moscow downlinked several hours worth of test data of the ASN-M Satellite Navigation System from ~9:05am-1:40pm via S-band.

SHERE Update: “Greg: wonderful work on SHERE yesterday! Once again, we’ve gotten more than we had hoped for, and thank you for the test repeats. Also, thank you for your observations and comments in the crew notes, especially on the uniformity and symmetry of the fluid column. It has started us thinking about the effects of fluid overflow on the Force Transducer end and optical effects in our camera field of view.”

No CEO (Crew Earth Observations) photo targets uplinked for today.

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

Week 19 Scheduled Main Activities:

  • Mon. (8/25): RPCM LAB1D5/AvRack R&R; O2 port install on WRS; ATV cargo transfers; SVO EDV condensate fill; TEPC relocate; Flowmeter battery maint.; ER5 rack transfer preps.
  • Tue. (8/26): SRVK KAV sample; ATV cargo transfers; ER4/ER5 stow; ER4 relocate to JPM; ER5 relocate to JPM; ER4/ER5 umbillicals mate; ER4/ER5 restow.
  • Wed. (8/27): MBI-12/Sonokard (FE-1); ATV cargo transfers; LULIN data dwnld.; DC1 IDZ-2 smoke detector cleaning; THC-IMV flow meas.; SVO BRP-M water sampling; IWIS setup; Emergency proc. OBT/drill; IWIS thruster test.
  • Thu. (8/28): MBI-12/Sonokard (CDR); BMP ch.1 regen; Regen ECLSS Mod Kit 1 install; JAXA CB Microscope C/O; Clay EPO return; COL FSL FCE release; ATV cargo transfers; PHS w/blood set-up; ASU R&R; WRM audit.
  • Fri. (8/29): BMP ch.2 regen; PHS w/blood (PCBA); 29P: LKT remove/activate/hatch close/leak check; BRTK-MBRL prox.comm. prep; ITCS fluid sampling (JEM, Lab, Node-2, COL).

ISS Orbit (as of this morning, 8:01am EDT [= epoch]):
Mean altitude — 355.7 km
Apogee height — 361.6 km
Perigee height — 349.7 km
Period — 91.65 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0008791
Solar Beta Angle — -30.2 deg (magnitude decreasing)
Orbits per 24-hr. day — 15.71
Mean altitude loss in the last 24 hours — 42 m
Revolutions since FGB/Zarya launch (Nov. 98) — 55928

Significant Events Ahead (all dates Eastern Time, some changes possible.):
08/25/08 — Progress M-64/29 SORR (Stage Operations Readiness Review) @ NASA/JSC
09/01/08 — Progress M-64/29P undocking, from FGB nadir; de-orbit/re-entry ~9/8
09/05/08 — ATV1 undocking, from SM aft port (loiter until 9/29 for nighttime reentry/observation)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking (SM aft)
10/01/08 — NASA 50 Years (official)
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft)
10/12/08 — Soyuz TMA-13/17S launch (~3:03am EDT; Lonchakov, Fincke, Garriott)
10/14/08 — Soyuz TMA-13/17S docking (FGB nadir port, ~4:51am)
10/23/08 — Soyuz TMA-12/16S undocking (DC1 nadir) or 10/24?
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/25/08 — Progress M-65/30P undocking & deorbit
11/26/08 — Progress M-66/31P launch
11/30/08 — Progress M-66/31P docking
02/09/09 — Progress M-66/31P undocking & deorbit
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment
02/14/09 — STS-119/Discovery/15A docking
02/24/09 — STS-119/Discovery/15A undocking
02/26/09 — STS-119/Discovery/15A landing (nominal)
03/25/09 — Soyuz TMA-14/18S launch
03/27/09 – Soyuz TMA-14/18S docking (DC1)
04/05/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress M-67/32P undocking & deorbit
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/25/09 — Soyuz TMA-15/19S launch
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking)
07/30/09 — STS-128/Atlantis/17A – MPLM(P), last crew rotation
10/15/09 — STS-129/Discovery/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P)
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 (contingency).

SpaceRef staff editor.